What is the activity of a 20.1 μCi sample of carbon-14 in becquerels?
1
Expert's answer
2014-11-26T11:26:36-0500
1 Ci = 3.7 × 1010 Bq
a = 3.7× 1010×20.1 = 7.44 × 1011 Bq